For a one-tailed hypothesis test (upper tail) the p-value is computed to be 0.034. If the test is being conducted at 95% confidence, the null hypothesis Select one: a. is rejected. b. is not rejected. c. could be rejected or not rejected depending on the value of the mean of the sample. d. could be rejected or not rejected depending on the sample size.
Answer #1

For one tailed Hypothesis test the p value is 0.034

Rejection Rule: Reject the null hypothesis (H0) at \alpha level of significance when p value is less than \alpha

Here test was conducted at 95% confidence

(1-\alpha)=.95

\alpha=1-.95=0.05

p value=0.034 is less than \alpha=0.05 hence we will reject the null hypothesis.

Answer is a The null hypothesis is rejected.